Can someone help me find the inverse Laplace of:

10/(s(s^2+2s+5))
Answer #1

het, 10 Y(s) = SCS²+25+5) We decompose this into partial fractions. Let , 10 Bs & C S2+2 3+5 J 3(52+25+5) 10 3 (32+2 5+5) So,- .: 2612+2.1+5) + 113+C) = 10 2 x 8 & (B+C) = 10 = B + 0 = - 6- A(1²_2+5)-(-B + C) = 10 i. 2 x 4 +(B-C) = 10 - B - C = 2 - P